The flow rate of blood through a 2.00x10-9-m -radius
capillary is 3.80x10 cm /s . (a) What is the speed of the
blood flow? (This small speed allows time for diffusion of
materials to and from the blood.) (b) Assuming all the blood in
the body passes through capillaries, how many of them must
there be to carry a total flow of 90.0 cm /s ? (The large
number obtained is an overestimate, but it is still reasonable.)
1
Expert's answer
2019-11-19T07:46:16-0500

a)


"q=Av=\pi r^2 v"

"3.8\cdot 10^{-9}=\pi (2\cdot 10^{-4})^2v"

"v=0.0302\frac{cm}{s}"

b)


"N=\frac{Q}{q}"

"N=\frac{90}{3.8\cdot 10^{-9}}=2.37\cdot 10^{10}"